OpenStudy (anonymous):

28- 5e = 4(10- 3e) + 6e

OpenStudy (anonymous):

12 = e

OpenStudy (anonymous):

thank u so much but do u mind showing the steps?

OpenStudy (anonymous):

28-5e=4(10-3e)+6e 28-5e=40-12e+6e 28-5e=40-6e 28=40+e 28-40=-12 therefore making -12=-e which makes 12=e.

OpenStudy (anonymous):

Do you understand what I did?

OpenStudy (anonymous):

thanks! :D and yes

OpenStudy (anonymous):

No problem :D